SAVINGS Mr. Harris saves $20,000 in a money-market account at an interest
rate of 5.2%.
a. Determine the value of his investment after 10 years if interest is
compounded quarterly.
b. Determine the value of his investment after 10 years if interest is
compounded continuously.

Answer by Boreal(15235)

P=Po20000(1+(.052/4)^40=$33528.01. Round at the end.
P=Poe^(.052*10)=20000e^(.52)=$33640.55